OpenStudy (anonymous):

Simplify the given expression to rational exponent form and justify each step by identifying the properties of rational exponents used. All work must be shown.

OpenStudy (anonymous):



OpenStudy (anonymous):

Can someone please help, I don't know how to do this

OpenStudy (anonymous):

@phi hate to bother you but youre the only math person that's on .-.

OpenStudy (phi):

\[ \frac{1}{\sqrt[3]{x^{-6}}} \]?

OpenStudy (anonymous):

Yes except its 3 not 2

OpenStudy (phi):

the trick with test problems is to change roots to *fractional exponents the cube root means the same thing as using an exponent of ⅓ * it's tiny font, but it says 3 (not 2). maybe you can "zoom in"

OpenStudy (anonymous):

zoomed in now, easier to read ok so root to fraction exponent but how do you do that?

OpenStudy (anonymous):

i'll look in lesson and see if i can find out

OpenStudy (phi):

get rid of the \( \sqrt[3]{ original} \) and write \[ ( original )^\frac{1}{3} \]

OpenStudy (phi):

It means the same thing. but we can use this rule \[ (a^b)^c = a^{bc} \]

OpenStudy (anonymous):

(x^-6)^1/3 so would it look like this?

OpenStudy (phi):

yes but remember that is all under 1 (don't forget)

OpenStudy (phi):

\[ \frac{1}{\left( x^{-6}\right)^\frac{1}{3}} \]